OpenStudy (selena2345):

In a school election the winning candidate got 44 votes.This was 55% of the total vote.How many students voted in the election ?

OpenStudy (anonymous):

80

OpenStudy (selena2345):

how can you explain

OpenStudy (anonymous):

cross multiply, 44/x = 55/100. so x = 80

OpenStudy (selena2345):

thank you